# ConsRec: Denoising Sequential Recommendation through User-Consistent Preference Modeling

### v1

@@ A.7 Theoretical Analysis of Consistent User Preference Modeling via ConsRec
Preliminary. Let =$\mathcal{H}=\{v_{1},\dots,v_{n}\}$ be the full interaction history of a user. The consistent user preference set $P$ can be represented:
$P\subseteq\mathcal{H},\lvert P\rvert=m,$ (17)
@@
$N=\mathcal{H}\setminus P,\lvert N\rvert=n-m.$ (18)
@@
$G_{\mathcal{U}}=(\mathcal{H},\mathcal{E}).$ (19)
@@
$C_{\text{max}}=\mathrm{arg}\max_{1\leq k\leq\ell}\lvert C_{k}\rvert$ (21)
@@
(i) Every pair of nodes belongs to $P$ is similar:
$\forall v_{i},v_{j}\in P:S_{ij}\geq\tau.$ (22)
@@
$\forall\{C_{1}^{N},\ldots,C_{q}^{N}\}\subseteq\mathcal{C}:\max_{1\leq r\leq q}\lvert C_{r}^{N}\rvert\mathrm{<}m.$ (24)
Lemmas. The following three lemmas are correspond to assumptions (i)-(iii), respectively.
(1) The preference set is the connected component. By assumption (i), $\forall v_{i},v_{j}\in P$, then $S_{ij}\geq\tau$, so that $(v_{i},v_{j})\in\mathcal{E}$. If any two nodes in the set $P$ are directly connected to each other, then $P$ forms a complete graph, which necessarily constitutes a connected component.

### v2

@@ A.3. Theoretical Analysis of Stable User Preference Modeling via LISRec
Preliminary. Let $\mathcal{H}=\{v_{1},\dots,v_{n}\}$ be the full interaction history of a user. The consistent user preference set $P$ can be represented:
(15) $\small P\subseteq\mathcal{H},\lvert P\rvert=m,$
@@
(16) $\small N=\mathcal{H}\setminus P,\lvert N\rvert=n-m.$
@@
(17) $\small G_{\mathcal{U}}=(\mathcal{H},\mathcal{E}).$
@@
(19) $\small C_{\text{max}}=\mathrm{arg}\max_{1\leq k\leq\ell}\lvert C_{k}\rvert$
@@
(i) The preference nodes are mutually reachable in $G_{\mathcal{U}}$:
(20) $\small\forall v_{i},v_{j}\in P:v_{i}\sim_{\tau}v_{j}$
@@
(22) $\small\forall\{C_{1}^{N},\ldots,C_{q}^{N}\}\subseteq\mathcal{C}:\max_{1\leq r\leq q}\lvert C_{r}^{N}\rvert\mathrm{<}m.$
Lemmas. The following three lemmas correspond to assumptions (i)-(iii), respectively.
(1) The preference set induces a connected component. By assumption (i), every pair of nodes in $P$ can reach each other through edges whose similarities exceed $\tau$, hence all nodes of $P$ lie in the same connected component of $G_{\mathcal{U}}$. Moreover, by assumption (ii), no node in $N$ is adjacent to any node in $P$, so that component contains only the nodes of $P$.